# Preprint version

### v1

@@ 1 Introduction
Deep learning models have enabled significant improvements in Time Series (TS) classification, due to their powerful representational capabilities [1]. However, most models suffer from a lack of generalization and can hardly be transferred from their training domain (the so-called source domain) to an unfamiliar, external domain (the target domain) [2]. This limitation is mainly caused by distribution shifts, a very common phenomenon in TS [3, 4]. To address this problem, Unsupervised Domain Adaptation (UDA) attempts to find a domain-invariant feature space for labeled source samples and unlabeled target samples. Recently, several UDA algorithms have been benchmarked for TS tasks [2], mainly using image processing architectures. The limited performance of UDA methods on TS datasets compared to their high accuracy on image tasks underscores the need for TS-specific architectures. Since [2], many approaches have enhanced the performance of TS representation by identifying invariant temporal features specifically tailored for TS [5, 6, 7, 8, 9]. For example, frequency-based features extracted from neural operators, such as Fourier Neural Operators (FNO) [10], have improved domain-invariant representations and have been used for UDA on TS [11]. The UDA framework actually covers a special case where the classes are the same in both domains (despite some small drift between domains) known as the closed-set assumption. Motivated by this limitation, Universal Domain Adaptation (UniDA) was introduced [12]. In UniDA, the target domain remains unlabeled, certain classes are shared between the source and target domains, while others are exclusive to either domain. Unlike standard UDA, UniDA helps to discover unknown class samples that are unique to the target domain. This capability is crucial in TS classification, where identifying unseen patterns is often critical [13].
@@
Few approaches attempt to address the dynamic nature of the threshold value. For example, TNT [25] attempts to compute a task-specific threshold, but this computation is based solely on the source dataset. Domain shifts may render such computations ineffective on the target dataset. Moreover, OVANet [26] employs a One-vs-All strategy that eliminates the need for a threshold to distinguish common samples from unknown ones [27]. In addition to the challenge of defining a threshold, CMU [24] states that relying on a single discriminability metric for OOD detection (such as the entropy used in UAN) is not sufficiently accurate and proposes a combination of metrics (entropy, consistency, and confidence). The lack of discriminability of a single metric often arises from the overconfidence exhibited by deep learning-based models [28], which limits their ability to generate uncertainty spaces for unknown samples.
Finally, the previously described UniDA methods were originally designed for computer vision tasks. To the best of our knowledge, the only UniDA approach tailored for TS is RAINCOAT [11], which introduces a modification of the FNO layer. In this method, an OT alignment is performed based on the variation of the target samples’ latent representations in a deep reconstruction scheme. This suggests that the presence of private class samples will modify the representation space after some optimization under a reconstruction loss. Such variation is then measured by a statistical test for bimodality on the target batch samples. The underlying RAINCOAT assumption is that: pseudo-labelling common and unknown classes is possible at the batch level. This approach is computationally expensive due to its reliance on this internal deep learning optimization step. However, some elements can be retained to approach UniDA on TS: the FNO architecture and its pseudo-labelling assumption.
@@ 2.1 Fourier Neural Operator
The Fourier Neural Operator (FNO) was introduced in [10] with a focus on approximating the solution operators of partial differential equations. The FNO relies on multiple Fourier layers, which consist of projecting the output of a neural network into the Fourier space by applying a Fast Fourier Transform. Following this step, a linear transformation is applied to the lower Fourier modes while the higher modes are filtered. Finally, an Inverse Fast Fourier Transform is applied to project back the data into the original space.
The original Fourier layer is slightly modified in [11], by using a cosine smoothing function to prevent alignment over noisy frequency features and extracting polar coordinates of the frequency coefficients instead of performing an inverse Fourier transform. In an abuse of notation, this modified Fourier layer will be referred to as FNO in this paper.
@@ 2.3 Domain Adaptation with DeepJDOT
Let $g$ be a feature extractor that maps the input space into a feature space, and $f$ be a classifier that maps the feature space to the label space. Let $h(x)=f(g(x))$. DeepJDOT [21] proposes to minimize the joint discrepancy between the distributions of each domain seen in both the feature and label spaces while optimizing the classification accuracy on the target domain.
This can be done by defining the following transportation cost for each source sample $x_{i}^{s}$ and target sample $x_{j}^{t}$:
@@
The optimization scheme follows an alternating approach: first set $\gamma$, then set $f$ and $g$. Setting $f$ and $g$ reduces the problem to an OT minimization (1), while setting $\gamma$ makes optimizing $f$ and $g$ a standard deep learning task.
DeepJDOT is a UDA method that assumes that the classes in both the source and target domains are the same. The alignment of the domains results in a small discrepancy between the source and target distributions at the end of the training process.
@@ 2.4 Universal Domain Adaptation
UniDA extends the scope of UDA by addressing a more general and challenging setting [12]. In UniDA, we are given a labeled source domain, $\mathcal{D}^{s}=\{(x_{i}^{s},y_{i}^{s})\}_{i=1}^{n_{s}},$ which follows a joint distribution $\mathcal{P}^{s}(x^{s},y^{s})$, and an unlabeled target domain, $\mathcal{D}^{t}=\{(x_{i}^{t})\}_{i=1}^{n_{t}},$ which follows another joint distribution $\mathcal{P}^{t}(x^{t},y^{t})$. As in UDA, it is assumed that the source and target distributions are not identical, i.e., $\mathcal{P}^{s}\neq\mathcal{P}^{t}$.
UniDA operates under the flexible assumption that the label sets may differ, i.e., $\mathcal{Y}^{s}\neq\mathcal{Y}^{t}$. To capture this distinction, the label sets are divided into three subsets [12]:
@@ 3.1 Pseudo-labelling
Given a pretrained network $h$ on the source domain, our goal is to discriminate between unknown and common target data based on the classifier’s outputs. However, the classifier often exhibits overconfidence when dealing with unknown samples [28]. To mitigate this, we propose a merged approach that regularizes the classifier’s predictions by integrating distance-based probabilities within the feature space, while leveraging all target samples in a given batch. We assume that at the batch level, samples of both common and unknown classes exist.
As illustrated in Figure 2, each probability vector $h(x_{t})$ of the batch is adjusted by a distance-based probability vector, resulting in a collection of probability vectors $p_{t}^{\prime}$. The maximum value of each probability vector in $p_{t}^{\prime}$ is then assembled into a histogram. Based on this histogram, an automatic thresholding approach separates the samples of the target batch into common and unknown samples.
More precisely, for a given sample $x_{t}$, we condition the classifier outputs $h(x_{t})$ with distance-based probabilities using the softmax function $\sigma$:
$p_{t}^{\prime}=h(x^{t})\sigma(-d_{t}),$
@@ 3.2 Anchors
Only target samples in $\mathcal{X}_{\mathcal{C}}^{t}$ should be aligned with source samples. Therefore, we need to introduce anchors to align target samples in $\mathcal{X}_{\mathcal{U}}^{t}$ to avoid misalignment with source samples. Since DeepJDOT relies on joint alignment in the feature space and the decision space, anchors should be defined in both spaces. On the one hand, due to the well-defined geometry of the decision space as a simplex, we choose $r=\frac{1}{|\mathcal{C}|}\mathbbm{1}$ as the only decision space anchor, since it corresponds to the uniform distribution over classes, representing maximum uncertainty and lowest confidence for any classifier. On the other hand, to define anchors in the feature space, we propose to use multiple anchors of feature vectors $\{a_{l}\}_{l=1}^{L}$ computed from unknown target samples. At initialization, a K-means algorithm computes the $L$ centroids of all the target sample feature vectors that form the anchors set. For each batch at the training stage, this set is updated using a moving average to reduce the computational complexity. The number $L$ takes into account that unknown samples may belong to multiple clusters in the feature space.
@@ 3.4 Inference
The pseudo-labelling module is inherently batch-dependent during training. However, batches as such do not exist during inference. Therefore, the pseudo-labelling assumption does not hold.
In order to address this limitation, we propose a method to set the threshold value after training based on the target training data used for adaptation. Specifically, after training, we can retain the last threshold computed by the model and use it during inference. Alternatively, we adopt a more robust approach that uses a larger validation batch and retains the automatically determined threshold.
@@ 4.1 Experimental Settings
UniDA performance is measured as in [18, 19, 26], using the H-score defined as $(2A_{\text{C}}A_{\text{U}})/(A_{\text{C}}+A_{\text{U}})$, where $A_{\text{C}}$ is the accuracy on target common classes and $A_{\text{U}}$ on target unknown classes. We compare UniJDOT with five UniDA state-of-the-art baselines: UDA, OVANet, DANCE, PPOT, and UniOT. It is important to mention that RAINCOAT’s reproducibility issues with the UniDA task prevented us from including it in the experiments. These problems were also reported by several researchers [^1] See: https://github.com/mims-harvard/Raincoat/issues/10.
@@ 4.2 Results
Scenario UDA OVANet PPOT DANCE UniOT UniJDOT
@@
Scenario UDA OVANet PPOT DANCE UniOT UniJDOT
@@
Scenario UDA OVANet PPOT DANCE UniOT UniJDOT
@@
Tables 1, 2 and 3 respectively report the H-scores of multiple adaptation scenarios for HAR, HHAR and EDF. Each scenario was tested over 10 different seeds. The error bars correspond to the standard error over these 10 seeds. The results are obtained using the best feature extractor (CNN or CNN+FNO) found during the hyperparameter search for each model. Note that for OT-based methods, the CNN+FNO architecture demonstrates better results. On average, our method consistently outperforms all baselines for all datasets. In most scenarios, our approach achieves first or second-best results.
While several methods, such as UniOT, PPOT, and UDA, show promising results on HAR, the performance gap between these methods and ours tends to widen on HHAR and EDF. Notably, the hyperparameters for all models, including ours, were chosen on randomly selected tasks from the HAR dataset. This observation suggests that these models have overfitted to HAR, resulting in reduced performance on other datasets such as HHAR and EDF.
Many UniDA models, including PPOT [18], UDA [12], and OVANet [26], rely on manually selected thresholds to distinguish between common and unknown target samples. This reliance on fine-tuned thresholds appears to be particularly detrimental to PPOT’s performance on HHAR, as the model struggles to converge effectively. In contrast, our proposed method, which avoids the use of a fixed or tunable threshold, demonstrates greater robustness across datasets. This adaptability allows our model to maintain superior performance even when evaluated on different datasets, highlighting its generalizability and effectiveness in UniDA scenarios.
(a) Interdataset threshold sensitivity
(b) Intradataset threshold sensitivity
Figure 4: Threshold sensitivity: The stars correspond for a) each dataset or b) each scenario. The dot lines show UniJDOT scores when using auto-thresholding.
Sensitivity to the threshold
@@
Finally, Table 4 compares several well-known auto-thresholding techniques for image binarization—specifically, Yen [36], Otsu[33], Triangle [37] and Li [34, 35]. These binarization methods implicitly rely on a hidden hyperparameter: the number of histogram bins. However, this choice is not critical, as the number of bins can be over-parameterized beyond the batch size to achieve finer granularity without negatively impacting performance. Table 4 shows that Yen consistently outperforms the others, indicating that this method is robust across all evaluated datasets compared to the alternatives. Note that Otsu’s is systematically second and Li’s is systematically third, illustrating the stability and consistency of automated thresholding techniques in the UniDA context.
@@
Ablation study
The ablation study in Table 5 evaluates three core components of our approach: the joint decision mechanism, the auto-thresholding, and the FNO layer. To simulate the absence of auto-thresholding, we replaced it with the best fixed threshold for each dataset, determined directly from the test data. Specifically, we chose the threshold corresponding to the peak performance for each dataset, as shown in Fig. 4(a). In addition, in the absence of the FNO layer, only CNN layers were used. Table 5 illustrates that removing the auto-thresholding— and replacing it with an oracle-fixed threshold—consistently degrades performance. When both the joint decision mechanism and auto-thresholding are removed, performance drops significantly across all datasets. As for the FNO layer, it outperforms the CNN-only architecture on HAR and EDF and achieves the second-best performance on HHAR. Conversely, removing both FNO and the joint decision mechanism results in the lowest overall scores. Overall, the joint decision mechanism consistently improves performance across all three datasets. These results demonstrate the contributions of this work, in particular the proposed joint decision mechanism and auto-thresholding for improving unknown class detection, as well as the benefits of using a TS–oriented architecture such as FNO.
@@
Auto-Thr. Joint Dec. FNO HAR HHAR EDF

### v2

@@ 1 Introduction
Deep learning models have enabled significant improvements in Time Series (TS) classification, due to their powerful representational capabilities [1]. However, most models suffer from a lack of generalization and can hardly be transferred from their training domain (the so-called source domain) to an unfamiliar, external domain (the target domain) [2]. This limitation is mainly caused by distribution shifts, a very common phenomenon in TS [3, 4]. To address this problem, Unsupervised Domain Adaptation (UDA) attempts to find a domain-invariant feature space for labeled source samples and unlabeled target samples. Recently, several UDA algorithms have been benchmarked for TS tasks [2], mainly using image processing architectures. The limited performance of UDA methods on TS datasets compared to their high accuracy on image tasks underscores the need for TS-specific architectures. Since [2], many approaches have enhanced the performance of TS representation by identifying invariant temporal features specifically tailored for TS [5, 6, 7, 8, 9]. For example, frequency-based features extracted from neural operators, such as Fourier Neural Operators (FNO) [10], have improved domain-invariant representations and have been used for UDA on TS [11].
The UDA framework actually covers a special case where the classes are the same in both domains (despite some small drift between domains) known as the closed-set assumption. Motivated by this limitation, Universal Domain Adaptation (UniDA) was introduced [12]. In UniDA, the target domain remains unlabeled, certain classes are shared between the source and target domains, while others are exclusive to either domain. Unlike standard UDA, UniDA helps to discover unknown class samples that are unique to the target domain. This capability is crucial in TS classification, where identifying unseen patterns is often critical [13].
@@
Few approaches attempt to address the dynamic nature of the threshold value. For example, TNT [25] computes a task-specific threshold, but this computation is based solely on the source dataset. Domain shifts may render such computations ineffective on the target dataset. Moreover, OVANet [26] employs a One-vs-All strategy that eliminates the need for a threshold to distinguish common from unknown samples [27]. In addition to the challenge of defining a threshold, CMU [24] states that relying on a single discriminability metric for OOD detection (such as the entropy used in UAN) is not sufficiently accurate and proposes a combination of metrics (entropy, consistency, and confidence). The lack of discriminability of a single metric often arises from the overconfidence exhibited by deep learning-based models [28], which limits their ability to generate uncertainty spaces for unknown samples
Finally, the previously described UniDA methods were originally designed for computer vision tasks. To the best of our knowledge, the only UniDA approach tailored for TS is RAINCOAT [11], which introduces a modification of the FNO layer. In this method, an OT alignment is performed based on the variation of the target samples’ latent representations in a deep reconstruction scheme. This suggests that the presence of private class samples will modify the representation space after some optimization under a reconstruction loss. Such variation is then measured by a statistical test for bimodality on the target batch samples. The underlying RAINCOAT assumption is that: pseudo-labelling common and unknown classes is possible at the batch level. This approach is computationally expensive due to its reliance on this internal deep learning optimization step. However, some elements can be retained to approach UniDA on TS: the FNO architecture and its pseudo-labelling assumption. Using FNO allows us to build a UniDA approach tailored for TS, as in [11].
@@ 2.1 Fourier Neural Operator
The Fourier Neural Operator (FNO) was introduced in [10] with a focus on approximating the solution operators of partial differential equations. The FNO relies on multiple Fourier layers, which consist of projecting the output of a neural network into the Fourier space by applying a Fast Fourier Transform. Following this step, a linear transformation is applied to the lower Fourier modes while the higher modes are filtered. Finally, an Inverse Fast Fourier Transform is applied to project back the data into the original space. The original Fourier layer is slightly modified in [11], by using a cosine smoothing function to prevent alignment over noisy frequency features and extracting polar coordinates of the frequency coefficients instead of performing an inverse Fourier transform. In an abuse of notation, this modified Fourier layer will be referred to as FNO in this paper.
@@ 2.3 Domain Adaptation with DeepJDOT
Let $g$ be a feature extractor that maps the input space into a feature space, and $f$ be a classifier that maps the feature space to the label space. Let $h(x)=f(g(x))$. DeepJDOT [21] proposes to minimize the joint discrepancy between the distributions of each domain seen in both the feature and label spaces while optimizing the classification accuracy on the target domain. This can be done by defining the following transportation cost for each source sample $x_{i}^{s}$ and target sample $x_{j}^{t}$:
@@
The optimization scheme follows an alternating approach: first set $\gamma$, then set $f$ and $g$. Setting $f$ and $g$ reduces the problem to an OT minimization (1), while setting $\gamma$ makes optimizing $f$ and $g$ a standard deep learning task. Each problem is solved alternatively. DeepJDOT is a UDA method that assumes that the classes in both the source and target domains are the same. The alignment of the domains results in a small discrepancy between the source and target distributions at the end of the training process.
@@ 2.4 Universal Domain Adaptation
UniDA extends the scope of UDA by addressing a more general and challenging setting [12]. In UniDA, we are given a labeled source domain, $\mathcal{D}^{s}=\{(x_{i}^{s},y_{i}^{s})\}_{i=1}^{n_{s}},$ which follows a joint distribution $\mathcal{P}^{s}(x^{s},y^{s})$, and an unlabeled target domain, $\mathcal{D}^{t}=\{(x_{i}^{t})\}_{i=1}^{n_{t}},$ which follows another joint distribution $\mathcal{P}^{t}(x^{t},y^{t})$. As in UDA, it is assumed that the source and target distributions are not identical, i.e., $\mathcal{P}^{s}\neq\mathcal{P}^{t}$. UniDA operates under the flexible assumption that the label sets may differ, i.e., $\mathcal{Y}^{s}\neq\mathcal{Y}^{t}$. To capture this distinction, the label sets are divided into three subsets [12]:
@@ 3.1 Pseudo-labelling
Given a pretrained network $h$ on the source domain, our goal is to discriminate between unknown and common target data based on the classifier’s outputs. However, the classifier often exhibits overconfidence when dealing with unknown samples [28]. To mitigate this, we propose a merged approach that regularizes the classifier’s predictions by integrating distance-based probabilities within the feature space, while leveraging all target samples in a given batch. We assume that at the batch level, samples of both common and unknown classes exist. By doing so, we ensure that a relatively large distance in the feature space is associated with low model confidence.
As illustrated in Figure 2, each logit $h(x_{t})$ of the batch is adjusted by a distance-based probability vector, resulting in a collection of probability vectors $p_{t}^{\prime}$. The maximum value of each probability vector in $p_{t}^{\prime}$ is then assembled into a histogram. Based on this histogram, an automatic thresholding approach separates the samples of the target batch into common and unknown samples.
More precisely, for a given sample $x_{t}$, we condition the classifier outputs $h(x_{t})$ through element-wise multiplication with distance-based probabilities. The softmax function $\sigma$ is then applied to obtain a probability vector:
$p_{t}^{\prime}=\sigma\left(h(x^{t})\sigma(-d_{t})\right),$
@@ 3.2 Anchors
Only target samples in $\mathcal{X}_{\mathcal{C}}^{t}$ should be aligned with source samples. Therefore, we need to introduce anchors to align target samples in $\mathcal{X}_{\mathcal{U}}^{t}$ to avoid misalignment with source samples. Since DeepJDOT relies on joint alignment in the feature space and the decision space, anchors should be defined in both spaces. On the one hand, due to the well-defined structure of the decision space as a simplex, we choose $r=\frac{1}{|\mathcal{C}|}\mathbbm{1}$ as the only decision space anchor since it corresponds to the most uncertain class probability vector, representing lowest confidence for any classifier. On the other hand, to define anchors in the feature space, we propose to use multiple anchors of feature vectors $\{a_{l}\}_{l=1}^{L}$ computed from unknown target samples. At initialization, a K-means algorithm computes the $L$ centroids of all the target sample feature vectors that form the anchors set. For each batch at the training stage, this set is updated using a moving average to reduce the computational complexity. The number $L$ takes into account that unknown samples may belong to multiple clusters in the feature space.
@@ 3.4 Inference
The pseudo-labelling module is inherently batch-dependent during training. However, batches as such do not exist during inference. Therefore, the pseudo-labelling assumption does not hold. In order to address this limitation, we propose a method to set the threshold value after training based on the target training data used for adaptation. Specifically, after training, we can retain the last threshold computed by the model and use it during inference. Alternatively, we adopt a more robust approach that uses a larger validation batch and retains the automatically determined threshold.
@@ 4.1 Experimental Settings
UniDA performance is measured as in [18, 19, 26], using the H-score defined as $(2A_{\text{C}}A_{\text{U}})/(A_{\text{C}}+A_{\text{U}})$, where $A_{\text{C}}$ is the accuracy on target common classes and $A_{\text{U}}$ on target unknown classes. We compare UniJDOT with five UniDA state-of-the-art baselines: UAN, OVANet, DANCE, PPOT, and UniOT. It is important to mention that RAINCOAT’s reproducibility issues with the UniDA task prevented us from including it in the experiments. These problems were also reported by several researchers [^1] See: https://github.com/mims-harvard/Raincoat/issues/10.
@@
Scenario UAN OVANet PPOT DANCE UniOT UniJDOT
@@
Scenario UAN OVANet PPOT DANCE UniOT UniJDOT
@@
4.2 Results
Tables 1, 2 and 3 respectively report the H-scores of multiple adaptation scenarios for HAR, HHAR and EDF. Best results are reported in bold, and second-best results are underlined. Each scenario was tested over 10 different seeds. The error bars correspond to the standard error over these 10 seeds. The results are obtained using the best feature extractor (CNN or CNN+FNO) found during the hyperparameter search for each model. Note that for OT-based methods, the CNN+FNO architecture demonstrates better results. On average, our method consistently outperforms all baselines for all datasets. In most scenarios, our approach achieves first or second-best results.
While several methods, such as UniOT, PPOT, and UAN, show promising results on HAR, the performance gap between these methods and ours tends to widen on HHAR and EDF. Notably, the hyperparameters for all models, including ours, were chosen on randomly selected tasks from the HAR dataset. This observation suggests that these models have overfitted to HAR, resulting in reduced performance on other datasets such as HHAR and EDF.
Many UniDA models, including PPOT [18], UAN [12], and OVANet [26], rely on manually selected thresholds to distinguish between common and unknown target samples. This reliance on fine-tuned thresholds appears to be particularly detrimental to PPOT’s performance on HHAR, as the model struggles to converge effectively. In contrast, our proposed method, which avoids the use of a fixed or tunable threshold, demonstrates greater robustness across datasets. This adaptability allows our model to maintain superior performance even when evaluated on different datasets, highlighting its generalizability and effectiveness in UniDA scenarios.
(a) Interdataset threshold sensitivity
(b) Intradataset threshold sensitivity
Figure 4: Threshold sensitivity: The stars correspond for a) each dataset or b) each scenario. The dot lines show UniJDOT scores when using auto-thresholding. Each color is associated with a dataset.
@@ 4.2 Results
Scenario UAN OVANet PPOT DANCE UniOT UniJDOT
@@
4.2.1 Sensitivity to the threshold
@@
Finally, Table 4 compares several well-known auto-thresholding techniques for image binarization—specifically, Yen [36], Otsu[33], Triangle [37] and Li [34, 35]. These binarization methods implicitly rely on a hidden hyperparameter: the number of histogram bins. However, this choice is not critical, as the number of bins can be over-parameterized beyond the batch size to achieve finer granularity without negatively impacting performance. Table 4 shows that Yen consistently outperforms the others, indicating that this method is robust across all evaluated datasets compared to the alternatives. Note that Otsu’s is systematically second and Li’s third, illustrating the stability and consistency of automated thresholding techniques in the UniDA context.
@@
4.2.2 Ablation study
The ablation study in Table 5 evaluates three core components of our approach: the joint decision mechanism, the auto-thresholding, and the FNO layer. To simulate the absence of auto-thresholding, we replaced it with the best fixed threshold for each dataset, determined directly from the test data. Specifically, we chose the threshold corresponding to the peak performance for each dataset, as shown in Fig. 4(a). In addition, in the absence of the FNO layer, only CNN layers were used. Table 5 illustrates that removing the auto-thresholding and replacing it with an oracle-fixed threshold consistently degrades performance. When both the joint decision mechanism and auto-thresholding are removed, performance drops significantly across all datasets. As for the FNO layer, it outperforms the CNN-only architecture on HAR and EDF and achieves the second-best performance on HHAR. Conversely, removing both FNO and the joint decision mechanism results in the lowest overall scores. Overall, the joint decision mechanism consistently improves performance across all three datasets. These results demonstrate the contributions of this work, in particular the proposed joint decision mechanism and auto-thresholding for improving unknown class detection, as well as the benefits of using a TS–oriented architecture such as FNO.
@@ 4.2.2 Ablation study
Auto-Thresh. Joint Decision FNO HAR HHAR EDF
@@ 5 Conclusion
Acknowledgment
The authors acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE23-0004 (project ODD) and of the STIC AmSud project DD-AnDet under grant N°51743NC.